# Kashmir-190006, India

### v3

@@ 3.1 Hawking saddle
We proceed similarly for the left side bath and get these points in the corresponding $w^{\pm}$ coordinates. Note that the $w^{\pm}$ coordinates that we use to completely span the two-sided black hole, along with the two baths attached on either side. Then the Hawking saddle is the entanglement entropy between the two bath points using Cardy-Calabrese formulae.
@@
In Figure 5, the entropy curves corresponding to the old and new black holes are the Hawking saddle entropy curve for eternal black holes corresponding to the initial and final temperature, respectively. It can be seen that the entropy curve for our evaporating black hole starts as the old one but later becomes parallel to the new saddle. As expected, the Hawking saddle entropy curve becomes linear at late times, with slope proportional to the temperature of the black hole. We have now obtained the graph for our evaporating black hole, which shows a clear interpolation between the initial and final state.
3.2 Island saddle
$w(y_{1})$$b$$w(y_{2})$$a$$b$$a$
@@ 3.2 Island saddle
In the island saddle calculation, we iterate over all possible points ‘a’ (‘$y_{2}$’ in the above diagram 6) in a fixed time slice to find an extrema. Note that there is a symmetry between the left and the right side of the diagram.
@@
Figure 7 shows the decrease in island entropy over time from the black hole entropy corresponding to the black hole at the initial temperature to that corresponding to a black hole at the new temperature. The asymptotic behavior of the black hole coming to equilibrium in its final state is clearly reflected in the graph. As expected, the Page curve comes down for an evaporating black hole, whereas it stays at a constant value for an eternal black hole. Thus, we have directly obtained the Page curve after Page time for an evaporating black hole. Further, this numerical approach can be used to profile evaporation for any type of cooling that the bath might provide.
@@ 4.1 Bath Energy follows Newton’s law of cooling
In this case, we take $\frac{\pi c}{12\tilde{\beta_{t}}^{2}}=\frac{\pi c}{12}\left(\frac{1}{\tilde{\beta}^{2}}+\left(\frac{1}{\beta^{2}}-\frac{1}{\tilde{\beta}^{2}}\right)e^{-st}\right)$. We choose $s$ to be the cooling constant. We solve for $E$ and then solve for boundary parameterization. The rest of the machinery works similarly, replacing $\tilde{\beta}$ with $\tilde{\beta_{t}}$.
Finally, we want to see how the black hole entropy and the island entropy change with time for different values of $s$ and $k$. We mainly have three cases: $k>s,k<s$, and $k=s$. We also see how the energy of this black hole changes.
(a) Energy of black hole vs Boundary time
@@
(c) Entropy of Island saddle vs Boundary time
Figure 8: Case 1: $k>s$
Figure 8 shows how the energy, Hawking saddle entropy, and Island saddle entropy varies with time in the $k>s$ case. We see that the island entropy curve is slightly concave at early times and nearly flat at late times.
(a) Energy of black hole vs Boundary time
(b) Entropy of Hawking saddle vs Boundary time
(c) Entropy of Island saddle vs Boundary time
Figure 9: case 2: $k<s$
Figure 9 shows how the relevant quantities vary with time when $k<s$. We can see that the island entropy curve is nearly flat at early times and convex at late times.
(a) Energy of black hole vs Boundary time
(b) Entropy of Hawking saddle vs Boundary time
(c) Entropy of Island saddle vs Boundary time
Figure 10: Case 3: $k=s$
The behavior of the relevant quantites for $k>s$ case is summarised in Figure 10. We find that the island entropy curve is slightly concave at early times and slightly convex at late times.
4.2 Linear Decrease in Incoming Stress Tensor.
In this case, we take $\frac{\pi c}{12\tilde{\beta_{t}}^{2}}=\frac{\pi c}{12\beta^{2}}-st$. Here $s$ is the rate of decay in incoming energy. We follow the same procedure to calculate the island entropy for this evaporation profile. Here, the black hole energy as a function of time is initially concave and later becomes straight.
(a) Energy of black hole vs Boundary time
(b) Entropy of Hawking saddle vs Boundary time
(c) Entropy of Island saddle vs Boundary time
Figure 11: Linear decrease in bath stress tensor
Figure 11 shows the change in energy, Hawking saddle entropy, and Island saddle entropy with time for this case. We find the entropy curve to be slightly concave throughout. We do not expect the curve to be convex at later times, as there is no final temperature to settle in.
4.3 Quasi-static Evaporation
We would like to compare the results from the different models of evaporation to one where the cooling is quasi-static, with equilibrium being maintained at each time step. We model this by treating the black hole as an eternal black hole at temperature $\beta_{t}$ at each time step, with $\beta_{t}$ taken to be a constant during the timestep. The incoming and outgoing stress tensors are taken as the same function, thereby always ensuring equilibrium.
We can then vary $\beta_{t}$ as we like, and the entropy curve obtained will represent the island entropy for a quasi-static evolution. We take $\frac{1}{\tilde{\beta_{t}}^{2}}=\left(\frac{1}{\tilde{\beta}^{2}}+\left(\frac{1}{\beta^{2}}-\frac{1}{\tilde{\beta}^{2}}\right)e^{-kt}\right)$. This temperature profile corresponds to the temperature of a black hole evaporating due to a single sudden change in bath temperature thereby breaking breaking equilibrium. If we take this same temperature profile for a black hole evaporating in a quasi-static manner, it corresponds to cooling like Newton’s Law of Cooling for both the bath and the black hole.
Figure 12: Entropy of Hawking saddle vs Boundary time
Comparing Figure 12 with Figure 7, we see that the quasi-static evolution is a much faster decay. The difference in the numerical values of island entropy for the initial and final temperature is also much larger in magnitude. Note that this Figure 12 is for $k=0.01$, whereas for Figure 7, $k=0.1$. The value of $k$ directly correlates to the rate of change of temperature of the black hole. Even so, with a lower $k$, the drop in value for the island entropy for the quasi-static case is substantially more than that of a dynamically evaporating black hole.
As mentioned previously, the purpose of including this non-dynamical case is to provide a point of comparison for our previous results.
@@ 5 Conclusion
In this paper, we extended previous results on Page curves for evaporating black holes in JT gravity to the general case where the bath temperature can be controlled externally.
Using the Island formula, we tracked the Hawking and the Island saddle entropy numerically for an evaporating black hole. Unlike the previous discussions where a black hole was given a shock, and its evaporation back to its original state was studied, our approach is more general. We continuously change the bath temperature and consider several different profiles of continuous temperature change. In this paper, we studied three profiles: linear change, Newton’s law of cooling, and step function cooling. The result gives the Page curves of a slowly evaporating black hole in AdS under different external conditions of cooling, while remaining in the semiclassical regime.
Our numerical methods are entirely general and can then be used to study any cooling profile of the bath. We leave this to future investigations.
Physically, our result is expected to be useful in modeling black hole evaporation under general external conditions. Non-trivial environments can affect the evaporation of black holes. Our methods show how to track the evolution of the Page curve under such general conditions.

### v4

@@ 3.3.1 Hawking saddle
We proceed similarly for the left side bath and get these points in the corresponding $w^{\pm}$ coordinates. Note that the $w^{\pm}$ coordinates that we use to completely span the two-sided black hole, along with the two baths attached on either side. Then the Hawking saddle is the entanglement entropy between the two bath points using the Cardy-Calabrese formula (20).
@@
In Figure 5, the entropy curves corresponding to the old and new black holes are the Hawking saddle entropy curve for eternal black holes corresponding to the initial and final temperature, respectively. It can be seen that the entropy curve for our evaporating black hole starts as the old one but later becomes parallel to the new saddle. As expected, the Hawking saddle entropy curve becomes linear at late times, with slope proportional to the temperature of the black hole. We have now obtained the graph for our evaporating black hole, which shows a clear interpolation between the initial and final states.
3.3.2 Island saddle
[…]
@@ 3.3.2 Island saddle
In the island saddle calculation, we iterate over all possible points ‘a’ ($y_{2}$’ in Figure 6) in a fixed time slice to find an extrema. Note that there is a symmetry between the left and right sides of the diagram.
@@
Figure 7 shows the decrease in the island entropy over time from the black hole entropy corresponding to the black hole at the initial temperature to that corresponding to a black hole at the new temperature. The asymptotic behavior of the black hole coming to equilibrium in its final state is clearly reflected in the graph. As expected, the Page curve comes down for an evaporating black hole, whereas it stays at a constant value for an eternal black hole. Thus, we have directly obtained the Page curve after Page time for an evaporating black hole. In addition, this numerical approach can be used to profile evaporation for any type of cooling that the bath might provide.
@@ 4.1 Bath Energy follows Newton’s law of cooling
Newton’s law of cooling corresponds to the natural relaxation of a finite bath toward equilibrium with an external reservoir. It is widely used as an effective description in statistical mechanics.
In this case, we take $\frac{\pi c}{12\tilde{\beta_{t}}^{2}}=\frac{\pi c}{12}\left(\frac{1}{\tilde{\beta}^{2}}+\left(\frac{1}{\beta^{2}}-\frac{1}{\tilde{\beta}^{2}}\right)e^{-st}\right)$. We choose ‘$s$’ as the cooling constant. We first solve for ‘$E$’ and then solve for the boundary parametrization. The rest of the machinery works similarly, replacing $\tilde{\beta}$ with $\tilde{\beta_{t}}$.
We study how both the black hole entropy and the island entropy evolve for different relationships between the cooling constant $s$ and the parameter $k$. The three cases of interest are: $k>s,k<s$, and $k=s$. In each case, we also monitor the evolution of the black hole energy. For concreteness, we present numerical results for two representative choices: $s=k/2$ and $s=2k$.
(a) Black hole energy vs Time
(b) Island entropy vs Time
Figure 8: Energy and Entropy vs Time
In the figures 8(a) and 8(b) we see the change in energy of the black hole and its island saddle entropy, respectively. It appears that the Island entropy curve follows the energy curve. This is expected: the energy decreases, so does the area of the black hole. The island entropy decreases accordingly.
(a) Rate of change of energy
(b) Rate of change of island entropy
Figure 9: Rate of change of Energy and Entropy vs Time
Figure 10: Hawking Saddle Entropy
For Hawking radiation as well we can the slopes changing at different rates and finally converging to the slopes of the entropy curve for final state BH.
4.2 Linear Decrease in Incoming Stress Tensor
Linear decay of incoming stress tensor models scenarios in which the bath energy is drained at a constant rate. This provides a simple toy model that isolates the role of a time-dependent flux.
In this case, we take $\frac{\pi c}{12\tilde{\beta_{t}}^{2}}=\frac{\pi c}{12\beta^{2}}-st$. Here, $s$ is the rate of decay in the incoming energy. We follow the same procedure to calculate the island entropy for this evaporation profile.
Figures 8(a), 8(b), 10 show the change in energy, the Island saddle entropy, and the Hawking saddle entropy over time, respectively.
4.3 Quasi-static Evaporation
We would like to compare the results from the different models of evaporation with the quasi-static case. The quasi-static case is one where the equilibrium is maintained at each time step. We modeled this by treating the black hole as an eternal black hole at temperature $\beta_{t}$ at each time step, with $\beta_{t}$ taken as a constant during the timestep. The incoming and outgoing stress tensors are taken as the same function, thereby always ensuring equilibrium.
We can then vary $\beta_{t}$ as we like, and the entropy curve obtained will represent the island entropy for a quasi-static evolution. We take $\frac{1}{{\tilde{\beta}^{2}}_{t}}=\left(\frac{1}{\tilde{\beta}^{2}}+\left(\frac{1}{\beta^{2}}-\frac{1}{\tilde{\beta}^{2}}\right)e^{-kt}\right)$. This temperature profile corresponds to the temperature of a black hole evaporating due to a single sudden change in bath temperature, thereby breaking the equilibrium. If we use the same temperature profile for a black hole evaporating quasi-statically, the evolution corresponds to Newton’s law of cooling for both the bath and the black hole. Although we maintain the equilibrium with respect to 14, we further compute the boundary parametrization and the $w\pm$ coordinates numerically.
(a) Entropy of Island saddle vs Boundary time
@@ 4.3 Quasi-static Evaporation
Figure 11: Quasi static vs Instantaneous (delta function shock) change in bath temperature.
@@ 5 Conclusion
In this paper, we extended previous results on Page curves for evaporating black holes in JT gravity to more general cases where the bath temperature can be controlled externally.
We computed the Hawking and the Island saddle entropy numerically for an evaporating black hole in an environment with varying temperature. Unlike previous discussions in which a black hole was given a shock and its evaporation back to its original state was studied, our approach is more general. We continuously change the bath temperature and consider several different profiles of continuous temperature change. In this paper, we studied three profiles: linear change, Newton’s law of cooling, and step function cooling. The result gives the Page curves of a slowly evaporating black hole in AdS under different external conditions of cooling, while remaining in the semiclassical regime.
Our numerical methods are entirely general and can be used to study any cooling profile of the bath. Physically, our result is expected to be useful for modeling black hole evaporation under general external conditions.
An interesting extension of our analysis would be to study stochastic fluctuations in the bath temperature, modeling random or noisy environments. Since our method only requires a time-dependent specification of the incoming stress tensor, incorporating stochasticity is straightforward and would allow one to probe how robust the Page curve is to environmental noise.
Another natural generalization concerns coupling the black hole to multiple baths with competing temperatures. This would allow for the exploration of steady-state non-equilibrium situations where radiation continuously flows through the black hole system.
We plan to return to these problems in the future.